OpenStudy (anonymous):

Hi, does anyone knows how to integrate (sinx)/(2+x^2) dx from x=-1 to 1

OpenStudy (anonymous):

\[\int\limits_{-1}^{1} \frac{ sinx }{ (2+x^2) }dx\]

OpenStudy (kainui):

This integral is obviously 0, but do you want to really solve it or just get to the answeR?

OpenStudy (anonymous):

I have no idea how to start and pretty lost, so it'll be great ifyou can help me go through it steps by steps

OpenStudy (kainui):

And by obvious, I mean that since you have a sine function, it's an odd function, so any time you integrate it from some point that's symmetrical around 0, it's going to be 0. Since (2+x^2) is also symmetrical for all values since it's even, it will give the same weight at every point on the sine function.

OpenStudy (anonymous):

so even function cancel out odd function? as in sinx will cancel out the 2+x^2?

OpenStudy (kainui):

Yeah, so as for actually solving it, I really don't know how to do this one, and it would be painful to try! But once you understand that odd functions that are integrated from -a to +a is 0, it should make sense. Imagine adding up the area from the integral of x^3 or just plain sin(x) from some random negative distance to the same positive distance. You'll get an equal negative area on the left and an equal positive area on the right that cancel out.

OpenStudy (anonymous):

i'm going to try that now

OpenStudy (kainui):

The point of the even function on the bottom is that it won't change the left and right part's symmetry on the odd function. So now if you imagine cosine or x^2 then when you integrate from some random negative number to 0, it's equal to as if you had just integrated from 0 to that same number but positive now. So it won't affect how high and low the odd function is, so they'll still cancel out.

OpenStudy (kainui):

I'll be here doing some homework, think about it a bit and I'll check on you in a few minutes if you catch my drift! Good luck.

OpenStudy (anonymous):

thanks man! you're a legend, i'll just reread what you said til i understand the concept

OpenStudy (anonymous):

This one is hell to integrate by hand.. don't even know where to start

OpenStudy (kainui):

Try to draw a picture, or just calculate exact values of that integral between -1 and 1 at values like +1/2 and -1/2 to see that they will cancel out each other.

OpenStudy (anonymous):

Does anyone know how to actually calculate the indefinite integral for this question

OpenStudy (anonymous):

ahh i see it, i plugged in -1 and 1, since they're suppose to subtract each other, it will come out as 0!

OpenStudy (anonymous):

nope, i dont know how to do it with indefinite integral either, good luck man !

OpenStudy (kainui):

@ECE I CAN integrate this, but no sane person would integrate this perfectly symmetrical integral that equals 0 lol.

OpenStudy (kainui):

In case you're curious: http://www.wolframalpha.com/input/?i=sinx%2F(2%2Bx%5E2)dx&t=crmtb01

OpenStudy (anonymous):

yeah I put it in wolfram too lol

OpenStudy (anonymous):

thanks guys! I will close this question now

OpenStudy (anonymous):

Looks like you would need IBP then a substitution?

OpenStudy (anonymous):

Honestly no clue

OpenStudy (anonymous):

\[u = sinx, du = cosxdx\] \[dv = 1/2+x^2, v = \arctan(x/\sqrt{2})/\sqrt{2}\]

OpenStudy (anonymous):

Ah forget it, would take forever to solve lol

OpenStudy (alekos):

who comes up with these crazy questions?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

show that ∫(z)=z^3 is analytical,,,,,can any one help me here?